# Predictive Modeling of Power Outages during Extreme Events: Integrating Weather and Socio-Economic Factors

### v1

@@ 3.2 Input Data Balancing
A key challenge in predicting LPHC outages is the imbalance in outage data, where small-scale events dominate and large-scale extreme events are minority. This imbalance causes ML models to favor frequent low-impact cases, leading to poor performance on the critical but infrequent LPHC events. In order to balance the dataset, we apply SMOGN, which has been introduced to address the issue of imbalanced datasets in regression tasks [27]. It integrates two over-sampling technique, SMOTER and SMOTER-GN [28]. While the $L2norm$ filtering step extracts outages associated with extreme events, the SMOGN algorithm addresses the imbalance in outage severity within these events. SMOGN generates synthetic data by applying three key techniques: random under-sampling, SMOTER (Synthetic Minority Over-sampling Technique for Regression), and adding Gaussian noise. Random under-sampling involves randomly removing samples from the majority class. It generates new synthetic samples by interpolating rare cases. This is done by taking a weighted average of the target variable values between the original minority instance and its k-nearest neighbors [28].
@@
SMOGN is summarized in Algorithm 2: We mark a case as rare if its outage value is in the top 30% of the target distribution. A rarity threshold of 380 customers out was used and five nearest rare dataset neighbors were considered for SMOTER interpolation within a safe region. Also, Gaussian noise with perturbation 2% was added when neighbors lie outside the local safe range. Minority cases were oversampled approximately one-to-one, and the majority was undersampled to one half. Fig. 6 shows the influence of SMOGN on the outage occurrence probability distribution for the extreme weather dataset used in the case study.
The original dataset is dominated by small outages near zero, while the rebalanced dataset includes more high-consequence outages. After rebalancing, probability is reduced in the low-outage region and increased in the mid/high-outage region beyond the threshold.
@@ 3.3 Model Selection and Training
Support Vector Machine (SVM): SVM for regression identifies an optimal hyperplane in a higher-dimensional space that best fits the data while maximizing the distance from the nearest points. It minimizes errors, making it suitable for predicting continuous variables. SVM uses kernel functions to map the data into a space where these nonlinear patterns become easier to fit.
Adaptive Boosting (AdaBoost): It is a boosting algorithm for regression which strengthens weak models by prioritizing harder-to-predict instances, refining the ensemble’s accuracy with each iteration. The final prediction is a weighted sum of the predictions from all weak learners, with weights determined based on their performance. Given the imbalanced nature of outage data, AdaBoost can concentrate learning on rare and high-impact outage conditions.
@@ 3.4 Performance Evaluation
The following common performance metrics are used for models evaluation [29]:
Mean Absolute Error (MAE): It measures the average deviation of the predicted values from the actual observed values. It shows how far the model’s predictions are from the actual values on average.
[…]
Mean Square Error (MSE): That metric measures the average of the squared differences between the estimated values and the actual observed values. Lower MSE indicates that the model is more consistent and accurate across the full range of observations.
$MSE=\frac{1}{n}\sum_{i=1}^{n}(P_{o}(i)-\hat{P_{o}(i)})^{2}$ (8)
In both expressions, $P_{o}(i)$ denotes the actual observed value and $\hat{P_{o}(i)}$ represents the predicted value.
@@ 3.5 Outage Prediction Model Formulation
With all relevant features integrated into the dataset, the problem is formulated as predicting county-level power outages under extreme events. Eq. 3.5 shows the model formulation.
$P_{o}(t)$ $=f\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$ (9)
Here, $P_{o}(t)$ denotes the ratio of customers without power in a county at time $t$ during an extreme event, and $P_{o}(t-k)$ is the corresponding ratio observed during the $k$-th most recent extreme event. $W(t)$ is the vector of contemporaneous weather features for the county, while W(tk)W(t-k) contains the weather features recorded for the $k$-th most recent event. $S$ comprises socio-economic indicators, including average household income, unemployment rate, and the distribution of year built for residential structures. $I_{f}$ aggregates power-system attributes such as counts of poles, towers, substations, transformers, and line segments by county.
@@ 4 Experimental Setup and Simulation Results
The dataset is divided into training (2014 to 2018), validation (2019), and test (2020–2021) sets to evaluate the performance of the models. The implementation of RF, SVM, and AdaBoost is done using Python Scikit-learn library [30] and LSTM is implemented using Tensorflow library [31].
We conducted hyperparameter tuning utilizing the Grid- Search tool in Python [30], employing cross-validation to identify the optimal input parameters for the models, as shown in Table 1. The LSTM network is a one-layer LSTM with 128 nodes followed by a Dense layer. The LSTM layer employs its standard activation structure, where the cell state is updated using a tanh activation and the gating mechanisms use sigmoid functions. Adam optimizer is used with a learning rate of 0.001 and a batch size of 128.
Table 1: Cross-Validation Hyper-Parameters Selection
Hyper-Parameter RF SVM AdaBoost
Kernel rbf
Learning Rate - - 0.001
Estimators 4 - 120
Loss square
C - 10 -
Max Depth 100 - -
Gamma - 0.001 -
Fig. 7 illustrates the correlation analysis between socio-economic, infrastructure features and power outages. Higher mean income, lower unemployment rate, and more extensive infrastructure, including towers, lines, transformers, and poles, are associated with lower outages, suggesting that wealthier counties or those with more infrastructure might have higher power systems resilience. Substations and the year of structure construction also show a slight correlation, indicating newer or more substations may lead to improved outage resistance.
Figure 7: Pearson’s correlation of the customers outage number with the infrastructure and socio-economic variables.
@@
We also conducted a feature importance analysis using the RF-based Mean Decrease in Impurity (MDI). Fig. 8 depicts the contribution of each feature to the model’s predictive accuracy. The analysis showed that weather variables dominate outage prediction, with precipitation (19%), wind speed (13.5%), and surface pressure (13.1%) being the top contributors. Non-weather features, including infrastructure density (3.5%) and socioeconomic indicators such as average income (1.4%), also exhibit contribution to the prediction.
@@
Figure 8: Features importance across weather, infrastructure, and socioeconomic variables.
After these relationships are established, the models are trained using different combinations of input data, both with and without the inclusion of socioeconomic, infrastructure, and weather features, so that the influence of each data layer on predictive performance can be examined. By comparing model accuracy across these configurations, the added value of these additional features is evaluated, and their overall impact on the outage prediction process is determined.
Scenario-1 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n)\big)$
Scenario-2 $=\big(P_{o}(t-1),P_{o}(t-2),\cdots,P_{o}(t-n),$
$\quad W(t),W(t-1),\cdots,W(t-n),S,I_{f}\big)$
Scenario 1 uses weather data from the event day and the preceding day, together with outage information from the day before the extreme event and seasonal weather patterns. This configuration enables the models to utilize historical outage behavior along with weather conditions to estimate future outages. Scenario 2 expands upon this by incorporating all data used in Scenario 1 together with infrastructure and socioeconomic information.
Table 2: Performance Comparison of Models in Scenario-1 and Scenario-2.
Model Scenario-1 Scenario-2
MSE MAE MSE MAE
RF 0.0231 0.0898 0.0222 0.0827
SVM 0.0175 0.0796 0.0172 0.0774
AdaBoost 0.0210 0.0735 0.0204 0.0730
LSTM 0.0127 0.0749 0.0086 0.0432
Figure 9: Performance comparison between models in terms of MSE.
Table. 2 presents the performance of each model and shows that the scenario incorporating socioeconomic and infrastructure information produces lower prediction error. This outcome provides evidence that incorporating socioeconomic and infrastructure variables enhances the model’s predictive capability.
Fig. 10 illustrates the performance of the model by comparing the predicted outage values with the actual recorded values for LSTM model. Each point represents a data sample from the validation period, and the diagonal line indicates the ideal one to one correspondence between predicted and observed outages. As shown in the figure, the majority of the predictions follow the general upward trend of the reference line.
Figure 10: Scatter-plots of predicted versus actual customer outage for LSTM model.
Fig. 11 illustrates the comparison between the predicted and actual customer outage counts for the June 6–8, 2020 flood event in Wayne County, Michigan using four models: RF, SVM, AdaBoost, and LSTM. The time series plots show how well each model follows the temporal evolution and magnitude of the observed outages. Among the four models, the LSTM provides the closest match to the actual outage pattern. This indicates that the LSTM is able to effectively exploit temporal dependencies in the sequence of weather and related features during the event window. The Random Forest model also exhibits good performance. The AdaBoost model shows moderate agreement with the actual outages. The SVM model presents the weakest performance among the four.
(a) RF prediction vs. actual outages
(b) SVM prediction vs. actual outages
@@
Figure 11: Comparison of predicted and actual outage counts for the June 2020 flood event in Wayne County.
@@
Study Dataset Features Used Best Model Prediction Resolution
Lee et al. [12] EAGLE-I + NWS alerts Weather features only XGBoost
($R^{2}$ = 0.998, RMSE = 2115 customers)
Event-level maximum outage prediction;
No hourly or county-level resolution
Cruz et al. [11] EAGLE-I + NASA weather + Census data Weather + Socio-economic RF: MSE = 5.93e-05
MAE = 0.00102
RMSE = 0.00770
Next-day average outage prediction;
Not high-resolution
Our Work (Proposed) EAGLE-I + Open-Meteo + Census + Infrastructure Weather + Socio-economic + Infrastructure + Seasonal indicators LSTM
(MSE = 0.0086, MAE = 0.0432)
County-level, hourly outage prediction during extreme events
Given that this model is primarily trained on data from extreme weather and socio-economic features, it is well-suited to be used for planning to enhance the resilience of power systems. Planning can be done through simulating the system’s performance across a range of severe weather scenarios.
5 Conclusions
In this paper, we presented a novel machine learning–based approach for predicting power outages caused by LPHC events. The method integrated historical outage data from the EAGLE-I dataset, weather data from Open-Meteo, as well as infrastructure and socio-economic information to construct predictive models. We addressed the challenges of limited data availability and missing records for LPHC events by applying KNN and SMOGN methods. We showed the correlations between socio-economic and infrastructure features with extreme power outages. Several machine learning models, including RF, SVM, AdaBoost, and LSTM networks, were compared. The experimental evaluation on Michigan counties demonstrated that the LSTM model consistently outperformed RF, AdaBoost, and SVM in terms of lowest error. Our analysis indicated that weather-related variables—particularly precipitation and wind speed were the primary drivers of outages, while infrastructure and socio-economic factors exerted secondary influences. This model can be adapted to other regions by incorporating localized data, including region-specific weather patterns and infrastructure conditions as well as local social-economic information. Similarly, extending the model to handle other types of extreme events (e.g., wildfires and earthquakes) could be achieved by adding relevant features and retraining the model on new data.

### v2

@@ 3.2 Input Data Balancing
A key challenge in predicting HILP outages is the imbalance in outage data, where small-scale events dominate and large-scale extreme events are minority. This imbalance causes ML models to favor frequent low-impact cases, leading to poor performance on the critical but infrequent HILP events. In order to balance the dataset, we apply SMOGN, which has been introduced to address the issue of imbalanced datasets in regression tasks N20 . It integrates two over-sampling technique, SMOTER and SMOTER-GN Nina25 . This algorithm addresses the imbalance in outage severity within these events. SMOGN generates synthetic data by applying three key techniques: random under-sampling, SMOTER (Synthetic Minority Over-sampling Technique for Regression), and adding Gaussian noise. Random under-sampling involves randomly removing samples from the majority class. It generates new synthetic samples by interpolating rare cases. This is done by taking a weighted average of the target variable values between the original minority instance and its k-nearest neighbors Nina25 .
@@
SMOGN is summarized in Algorithm 2: We mark a case as rare if its outage value is in the top 30% of the target distribution. A rarity threshold of 380 customers out was used. five nearest rare dataset neighbors were considered for SMOTER interpolation within a safe region. Also, Gaussian noise with perturbation 2% was added when neighbors lie outside the local safe range. Minority cases were oversampled approximately one-to-one, and the majority was undersampled to one half. Fig. 6 shows the influence of SMOGN on the outage occurrence probability distribution for the extreme weather dataset used in the case study. The original dataset is dominated by small outages near zero, while the rebalanced dataset includes more high-consequence outages, leading to a smoother and more balanced distribution for training. After rebalancing, probability is reduced in the low-outage region and increased in the mid/high-outage region beyond the threshold.
@@ 3.3 Model Selection and Training
Graph Neural Network (GNN): GNNs are deep learning architectures designed to operate on graph-structured data LiaoMPCE . In this study, counties are represented as nodes in a spatio-temporal graph, with edges connecting geographically proximate counties (spatial edges) and linking the same county across consecutive time steps (temporal edges). Each node is characterized by weather, socio-economic, and infrastructure features. A Graph Attention Network (GAT) architecture is employed to perform message passing, where the attention mechanism adaptively weights the influence of neighboring nodes, enabling the model to focus on the most relevant spatial-temporal relationships. The model follows the procedure outlined in Algorithm 3.
Algorithm 3 Spatio-Temporal Outage Prediction Using Graph Attention Network (GAT)
1:Input: $D\leftarrow$ Dataset with hourly weather, socio-economic, spatial and extreme events data
2:Target: $y\leftarrow$ Number of customers outage during extreme events
3:Step 1: Preprocessing
4:Normalize features $\mathcal{F}$, target $y$ using Min-Max scaling
5:Step 2: Node define
6:Assign node IDs based on (county, time)
7:Step 3: Graph Construction
8:Spatial edges: For each timestamp $t$, connect counties within 50 miles
9:Temporal edges: For each county $c$, link consecutive timestamps $t_{i}\rightarrow t_{i+1}$
10:Step 5: Two Layers GAT Model
11:Layer 1: 6464-dim GAT with 44 heads + ReLU
12:Layer 2: 3232-dim GAT + ReLU
13:Output: Linear layer for scalar prediction
14:Step 6: Model Training
15:for epoch =1=1 to 150150 do
16:  Forward pass: $\hat{y}=\texttt{GAT}(G)$
17:end for
18:Step 7: Post-processing
19:Predict scaled output $\hat{y}_{\text{scaled}}$
20:Inverse-transform to obtain actual outage: ${y}=\texttt{InverseScaler}(\hat{y}_{\text{scaled}})$
21:Output: Predicted customer outages $y$
Adaptive Boosting (AdaBoost): It is a boosting algorithm for regression which strengthens weak models by prioritizing harder-to-predict instances, refining the ensemble’s accuracy with each iteration. The final prediction is a weighted sum of the predictions from all weak learners, with weights determined based on their performance.
@@ 3.4 Performance Evaluation
The following common performance metrics for regression tasksplevris2022investigation are used for models evaluation:
Mean Absolute Percentage Error (MAPE): It quantifies the average absolute deviation between predictions and observations as a percentage of the observation.
$\mathrm{MAPE}(\%)=\left(\frac{1}{n}\sum_{i=1}^{n}\left|\frac{P(i)-\hat{P}(i)}{P(i)}\right|\quad\right)\times 100$ (10)
Lower values indicate better performance. For example, a MAPE of 8% implies an average error equal to 8%.
$R^{2}$ Score (Coefficient of Determination): It quantifies the proportion of variance in the observed data that is explained by the model’s predictions. An $R^{2}$ value of 100% indicates perfect prediction, while a value of 0 implies that the model performs no better than predicting the mean of the observed values.
$R^{2}=\left(1-\frac{\sum_{i=1}^{n}\left(P(i)-\hat{P}(i)\right)^{2}}{\sum_{i=1}^{n}\left(P(i)-\bar{P}\right)^{2}}\right)\times 100$ (11)
where $P(i)$ is the actual outage value at instance $i$, $\hat{P}(i)$ is the predicted value, $\bar{P}$ is the mean of the actual values, and $n$ is the number of instances.
@@ 3.5 Outage Prediction Model Formulation
With all relevant features integrated into the dataset, the problem is formulated as predicting county-level power outages under extreme weather events. Let $y(c,t)$ denote the number of customers without power in county $c$ at time $t$ during an identified extreme-weather period. Eq. 12 shows the model formulation.
$y(c,t)=f_{\theta}\!\left(\mathbf{Y}_{c,t}^{(n)},\;\mathbf{W}_{c,t}^{(n)},\;\mathbf{S}_{c},\;\mathbf{I}_{c}\right),$ (12)
where $f_{\theta}(\cdot)$ denotes a data-driven mapping with parameters $\theta$ learned from historical data.
The temporal outage history is represented as
$\mathbf{Y}_{c,t}^{(n)}=\big[y(c,t-1),\;y(c,t-2),\;\ldots,\;y(c,t-n)\big],$
and the corresponding historical weather features are given by
$\mathbf{W}_{c,t}^{(n)}=\big[\mathbf{W}_{c,t},\;\mathbf{W}_{c,t-1},\;\ldots,\;\mathbf{W}_{c,t-n}\big],$
where $\mathbf{W}_{c,t}$ denotes the vector of contemporaneous weather variables.
The vector $\mathbf{S}_{c}$ contains county-level socio-economic indicators, including average household income, unemployment rate, and residential building age distribution, while $\mathbf{I}_{c}$ represents power infrastructure attributes.
This formulation enables the predictive models to capture nonlinear temporal dependencies in outage evolution, the influence of extreme weather dynamics, and the modulating effects of infrastructure and socio-economic vulnerability on outage severity.
@@ 4 Experimental Setup and Simulation Results
The dataset spanning from 2014 to 2024 is used to train the model, while the June 6–8, 2020 flood event in Wayne County is selected for testing the performance. The implementation of RF and AdaBoost is done using Python Scikit-learn library N22 and LSTM is implemented using Tensorflow library N21 . The GNN model employed in this study is based on the GAT architecture using PyTorch Geometric.
Hyperparameter tuning was performed using the GridSearch tool in Python N22 , with cross-validation employed to determine the optimal input parameters for each model, as summarized in Table 1.
Table 1: Cross-Validation Hyper-Parameters Selection.
Hyper-Parameter RF AdaBoost GNN LSTM
Learning Rate - 0.001 0.01 0.001
Estimators 100 120 - -
Hidden Units - - 128 $\rightarrow$ 32 128
Activation - - ReLU sigmoid
Optimizer - - Adam Adam
Max Depth - - - -
Epochs - - 300 100
@@
Fig. 7 depicts the contribution of each feature to the model’s predictive accuracy. The analysis result showed that weather features, such as precipitation and average wind speed, have the most significant contribution to the model’s performance. The analysis confirms that weather variables dominate outage prediction, with precipitation (19%), wind speed (13.5%), and surface pressure (13.1%) being the top contributors. Non-weather features, including infrastructure density (3.5%) and socioeconomic indicators such as average income (1.4%), also exhibit contribution to the prediction.
@@
Figure 7: RF features importance across weather, infrastructure, and socioeconomic variables.
(a) Random Forest prediction vs. actual outages
(b) GNN prediction vs. actual outages
@@
Figure 8: Comparison of predicted and actual outage counts for flood event in Wayne County (June 6–8, 2020).
Fig. 8 illustrates the comparison between the predicted and actual customer outage counts for the June 6–8, 2020 flood event in Wayne County, Michigan using four models: RF, GNN, AdaBoost, and LSTM. The time series plots show how well each model follows the temporal evolution and magnitude of the observed outages. Among the four models, the LSTM provides the closest match to the actual outage pattern. This indicates that the LSTM is able to effectively exploit temporal dependencies in the sequence of weather and related features during the event window. The Random Forest model also exhibits good performance. The GNN model shows moderate agreement with the actual outages. The AdaBoost model presents the weakest performance among the four.
Fig. 9 displays the performance of each model. Fig. 9 confirming that LSTM (highest $R^{2}$) offers the most accurate predictions for outage events in this scenario. The RF and GNN model presents a promising deep learning-based approach with potential for further enhancement, especially for spatial-temporal generalization.
Figure 9: Performance comparison between models in terms of Error (MAPE) and Accuracy ($R^{2}$).
@@
Study Dataset Features Used Models Applied Prediction Setup and Results
Lee et al. N41 EAGLE-I outage data Historical weather RF, KNN, XGBoost. XGBoost achieved the best performance. The study forecasts state-level maximum outages within 12 hours of a weather alert; however, it does not offer high-resolution predictions of outages caused by extreme events.
Cruz et al. cruz EAGLE-I + NASA (weather), Census Bureau (socio-economic), FEMA Risk Index Weather + Socio-economic Regression, RF, XGBoost, KNN. RF outperform other models. The model predicts next-day average outages during Hurricane Matthew (Oct 2016). It does not offer high-resolution predictions.
Our Work (Proposed) EAGLE-I + Open-Meteo (weather)+ Census (socio-economic), Infrastructure data Weather + Socio-economic + Infrastructure + Seasonal indicators RF, AdaBoost, LSTM, GNN. LSTM is the best performer. The model achieves county-level validation with hourly resolution in predicting affected customers during the June 2020 Wayne County flood.
Given that this model is primarily trained on data from extreme weather and socio-economic features, it is well-suited to be used for planning to enhance the resilience of power systems. Planning can be done through simulating the system’s performance across a range of severe weather scenarios. By running these simulations, planners can assess vulnerabilities, analyze system responses, and make data-driven decisions on where to allocate resources effectively to enhance grid resilience under severe conditions.
5 Conclusion
In this paper, we presented a novel machine learning–based approach for predicting power outages caused by HILP events. The proposed method integrated historical outage data from the EAGLE-I dataset, weather data from Open-Meteo, as well as infrastructure and socio-economic information to construct predictive models. We addressed the challenges of limited data availability and missing records for HILP events by applying KNN and SMOGN methods. Several machine learning models, including RF, GNN, AdaBoost, and LSTM networks, were compared. These models were trained and tested across scenarios with varying complexity.
The experimental evaluation on 20 Michigan counties demonstrated that the LSTM model consistently outperformed RF, GNN and AdaBoost. Our analysis indicates that weather-related variables—particularly precipitation, wind speed, and surface pressure—are the primary drivers of outages, while infrastructure and socio-economic factors exert secondary influences.
The model can be adapted to other regions by incorporating localized data, including region-specific weather patterns and infrastructure conditions and local social-economic information. Similarly, extending the model to handle other types of extreme events (e.g., wildfires and storms) could be achieved by adding relevant features and retraining the model on new data.
Acknowledgment
This paper is supported by the Department of Energy, Solar Energy Technologies Office (SETO) Renewables Advancing Community Energy Resilience (RACER) program under Award Number DE-EE0010413. The authors acknowledge early-stage discussions related to the doctoral dissertation research of Nina Fatehi. The content of this material represents the views of the authors only and is not endorsed by or necessarily reflective of the views of the Department of Energy.